OpenStudy (anonymous):

Write a word problem for the equation 3x+13=4x+4

OpenStudy (ryan123345):

Hey there! I see that you are new to openstudy. This is a great website on getting help with questions of any subject. It is a fun way of learning! I wish you good luck on openstudy and dont forget to read the http://openstudy.com/code-of-conduct so you know the rules and regulations :) Hope you have fun!

OpenStudy (anonymous):

a) You and 2 friends are going to see a play, the total cost is $24, how much does each ticket cost?

OpenStudy (anonymous):

Thank you.

OpenStudy (anonymous):

You're welcome (:

OpenStudy (anonymous):

(a) John bought 3 discount CD's at the local music store. He then went to the check out counter to pay for his CD's. The total of the CD's was $ 24. How much was one CD ? 3x = 24 (divide each side by 3 to find x (the cost of each CD) 3x/3 = 24/3 x = 8 Each CD cost $8.

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

(b) Tom and his brother, Larry, went to the book store. They each had the same amount of money to spend.There was a book bin with several books marked the same price.Tom bought 3 books and $13 worth of pencils. Larry bought 4 books and $ 4 worth of pencils. How much did the books cost ? 3x + 13 = 4x + 4 (x being the cost of the books) 3x - 4x = 4 - 13 -x = - 9 x = 9 The books were $9 a piece

OpenStudy (anonymous):

Thank you :)

OpenStudy (anonymous):

not a problem...glad I could help....that second one was a bit trickier...lol :)